OpenStudy (anonymous):

In a quadrilateral MNOP the length of side MN is (3x - 1) units and the length of side ON is (x + 7) units. What value of x will make it a rhombus? 3 4 6 8

OpenStudy (aroub):

3x-1=x+7 3x-x=1+7 2x=8 x=4